OpenStudy (gabylovesyou):

The formula for degrees Celsius is: C = 5/9 (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 8 degrees Celsius is.

OpenStudy (anonymous):

9c=5(f-32).......multiplying both sides with 9 9/5c= f-32....... dividing both sides by 5 95c+32=f ............adding 32 on both sides @Gabylovesyou

OpenStudy (gabylovesyou):

what do i put? as my answer ?? What is part 1 and part 2?

OpenStudy (anonymous):

this is answer of part 1 9/5c+32=f

OpenStudy (anonymous):

now put c=8 in equation of part 1 you will have answer of part 2

OpenStudy (gabylovesyou):

what is the anwer for part 2

OpenStudy (neo92):

part 2 is bit confusing.. do u need 8 degrees Fahrenheit in degrees Celsius or vice versa?

OpenStudy (anonymous):

calculate it dear \[9\times8\div5+32\]

OpenStudy (gabylovesyou):

16.4??

OpenStudy (gabylovesyou):

@theyatin

OpenStudy (anonymous):

its (72/5) +32

OpenStudy (gabylovesyou):

46.4 ??

OpenStudy (anonymous):

yup exactly sweet

OpenStudy (gabylovesyou):

is this a good answer for the question??? 9/5c+32=f 9 x 8 / 5 + 32 72 / 5 + 32 14.4 + 32 46.4

OpenStudy (anonymous):

exactly dearo

OpenStudy (gabylovesyou):

k thank u1 ")